Calculate the magnitude of the force required to just move a 140g object along a horizontal surface if the coefficient of friction is 1.9
1
Expert's answer
2020-11-23T10:26:27-0500

Applying the Newton’s Second Law of Motion, we get (at the moment when the object begins to move the acceleration is zero):


"\sum F_x=ma_x=0,""F_{req}-F_{fr}=0,""F_{req}=F_{fr}=\mu N = \mu mg,""F_{req}= \mu mg=1.9\cdot 0.14\ kg\cdot 9.8\ \dfrac{m}{s^2}=2.61\ N."

Answer:

"F_{req}=2.61\ N."